### v2

# Leveraging chaotic transients in the training of artificial neural networks

Abstract
Traditional algorithms to optimize artificial neural networks when confronted with a supervised learning task are usually exploitation-type relaxational dynamics such as gradient descent (GD). Here, we explore the dynamics of the neural network trajectory along training for unconventionally large learning rates. We show that for a region of values of the learning rate, the GD optimization shifts away from purely exploitation-like algorithm into a regime of exploration-exploitation balance, as the neural network is still capable of learning but the trajectory shows sensitive dependence on initial conditions –as characterized by positive network maximum Lyapunov exponent–. Interestingly, the characteristic training time required to reach an acceptable accuracy in the test set reaches a minimum precisely in such learning rate region, further suggesting that one can accelerate the training of artificial neural networks by locating at the onset of chaos. Our results –initially illustrated for the MNIST classification task– qualitatively hold for a range of supervised learning tasks, learning architectures (including both shallow and deep multilayer perceptrons and convolutional neural networks) and other hyperparameters (different activation functions and weight regularisation), and showcase the emergent, constructive role of transient chaotic dynamics in the training of artificial neural networks.
The multilayer perceptron (MLP) is an archetypical method in supervised machine learning [1, 2], used to infer (regress or classify) complex input-output representations $x\to y$, where $x\in\mathbb{R}^{m}$ and $y$ is usually another vector for regression tasks or an element from a discrete set in classification tasks. Accordingly, an MLP is mathematically an overparametrized nonlinear function $\mathscr{F}(x;{\Omega})$, where $\Omega=\{w_{k}\}$ is the set of trainable parameters. Often represented as an (artificial neural) network [1, 2], the MLP is visualized as a mathematical graph composed of various stacked layers of interconnected nodes, where the first layer represents the input vector $x\in\mathbb{R}^{m}$. The edges connecting nodes in adjacent layers represent (parametric) affine transformations of the vector elements of one layer to the next one (the coefficients of such transformations, usually called the edge weights, belong to the parameter set $\Omega$), and the nodes integrate the incident linear compositions in the edges nonlinearly via what is called an activation function. The information thus flows from the input layer $x$ onwards until reaching the final, output layer whose nodes represent the elements of the output $y=\mathscr{F}(x;{\Omega})$.
Optimizing the set of parameters $\Omega$ so that the mismatch between $\mathscr{F}(x;{\Omega})$ and $y$ –the so-called loss function $\mathscr{L}(x;\Omega)$– is minimized is called the training process. In practice, training the MLP is usually done iteratively, being gradient descent (GD) in parameter space the gold standard, where
$\omega_{k}(t+1)=\omega_{k}(t)-\eta\partial_{\omega_{k}}\mathcal{L}(x;\Omega(t)),\ \forall\omega_{k}\in\Omega$ (1)
where $\eta$ is the so-called learning rate and $\partial_{x}:=\partial/\partial x$.
Observe that, along training, the MLP is effectively performing a trajectory in the (graph) space spanned by the set of parameters $\Omega$. We contend that reinterpreting the training process as a (high-dimensional, latent) graph dynamics [3] allows to inquire the inner workings of learning algorithms following physics-inspired and complexity-based epistemics [4, 5, 6]. Instead of tracking the scalar projection of the full dynamics given by the time evolution of the loss function $\mathscr{L}(x;\Omega(t))$, let us consider the actual network trajectory [7]. We argue that this change of focus –which amounts to tracking full graph trajectories of the form $(\Omega(0),\Omega(1),\Omega(2)\dots)$, where the set $\Omega(t)=\{\omega_{1}(t),\dots,\omega_{m}(t)\}$ incorporates the updated values of the so-called weights and biases of the neural architecture at training epoch tt– can provide valuable insights if at the same time we leverage principles and tools from network science [8, 9, 10, 11, 12, 13, 14] and dynamical systems [15, 16].
[…]
Our contention in this work is that, indeed, such transition takes place, and the emergence of exploration-like dynamics, induced by the onset of sensitive dependence on initial conditions –the hallmark of chaotic dynamics– in the dynamics of Eq. 1 yields a sweet spot where the balance of exploitation and exploration dynamics has a constructive, beneficial role for neural network learning. Let us clarify at this point that in this work we are neither considering intrinsically dynamic neural networks whose neurons show chaotic behavior [34, 35, 36, 37, 38, 39] nor considering the stability of the input-to-output trajectory in deep networks [40], instead we are considering the training (optimization) dynamics in a neural network whose output is not dynamic. Moreover, we argue that (i) for an often unexplored region of large-but-not-too-large values of $\eta$, the map is optimally interpolating exploitation and exploration search strategies, and (ii) that the average training time required for the network to learn a representation capable of generalizing –i.e. capable to reach a good performance on the test set– $\texttt{Test}=\{x_{i},y_{i}\}_{i=1}^{N_{\text{test}}}$ is minimized in such sweet spot. Interestingly, this coincides with a similar type of optimality emerging when the loss function’s Hessian asymptotically evolves over training towards its so-called edge-of-stability, i.e. when its largest eigenvalue approaches $2/\eta$ [41]. In a nutshell, we argue that the transition to an exploitation-exploration balance is achieved by leveraging the onset of transient chaotic mixing [20], and at such transition the system efficiently minimizes training time, resulting in a possible demonstration of Langton’s hypothesis [42].
Figure 1: Training loss trajectory of a neural network on the MNIST dataset for three different learning rates: $\eta=0.01$, $\eta=7.5$ and $\eta=20$.
To actually demonstrate our hypothesis, we initially consider a vanilla classification task: MNIST image classification [43]. For parsimony, we initially choose a shallow MLP (one hidden layer with 64 neurons) and a $\tanh$ activation function (the results for other supervised learning tasks, activation functions deeper architectures and other architectures with inductive biases are discussed in the Appendix B Figs. 813). After train/test split, this MLP is trained on a set $\texttt{Train}=\{(x_{i},y_{i})\}_{i=1}^{N}$ of $N={6\cdot 10^{4}}$ labelled handwritten images ($y_{i}$ is the label of the $i$-th image), and we use a cross-entropy loss function
$\mathcal{L}(x;\Omega)=-\frac{1}{N}\sum_{i=1}^{N}y_{i}\log\mathscr{F}(x_{i};\Omega).$ (2)
Training takes place by using Eq. 1. For the sake of simplicity, no regularization is initially added to the loss function (results with $L_{2}$ weight regularization are discussed in Appendix B Fig. 11). Additionally, we use traditional GD schemes (no mini-batch or SGD) and discard dropout, to remove any source of stochasticity to the network dynamics [18]. Finally, to assess the performance for different learning rates, a constant learning rate $\eta$ is fixed throughout training.
Fig. 1 plots the time evolution of the (training) loss function $\mathcal{L}(x;\Omega(t))\equiv{\cal L}(t)$ for three different learning rates $\eta$. Interestingly, such loss is only monotonically decreasing for the standard range of small values of the learning rate. For larger values of $\eta$ other dynamical behaviors are found: converging loss functions with non-monotonic, irregular transients and other dynamical attractors for extremely large $\eta$, where the MLP does not seem to make any useful learning.
To better characterize what particular change in dynamical behaviors in the network evolution is inducing these projections in the loss function, we now resort to recently-introduced graph-theoretical extensions of the Maximum Lyapunov Exponent [44, 45], designed to estimate sensitivity to initial conditions in network trajectories. The procedure consists in three steps: (i) for a fixed learning rate, we define a set of $q$ different network initializations ${\cal S}=\{\Omega(0)\}$. (ii) Around each concrete network initialization $\Omega(0)$, we build an $\epsilon$-ball formed by a set of $M$ small network ‘perturbations’ of ${\cal B}=\{\Omega(0)^{(j)}\}_{j=1}^{M}$. To do that, we perturb each trainable parameter $w_{k}\in\Omega$ with uniform noise $w_{k}^{\prime}=w_{k}+\xi$, $\xi\sim\textsc{Uniform}(-\epsilon,\epsilon)$. Defining the distance between two network initializations $\Omega,\Omega^{\prime}$ as the $L_{1}$ norm $d(\Omega,\Omega^{\prime})=\sum_{\omega_{k}\in\Omega}|\omega_{k}-\omega_{k}^{\prime}|$, it follows that each perturbed network $\Omega(0)^{(j)}$ is at most at distance $\textsc{Card}(\Omega)\cdot\epsilon$ from $\Omega(0)$. (iii) Then, following [44] we measure the expansion rate of closeby network trajectories by adequately averaging the divergence of the $M$ elements inside each $\epsilon$-ball throughout the action of the training dynamics in Eq. 1:
$\Lambda_{\Omega(0)}=\frac{1}{\tau}\ln\frac{M^{-1}\sum_{j=1}^{M}d_{j}(\tau)}{M^{-1}\sum_{j=1}^{M}d_{j}(0)},$ (3)
where $d_{j}(\tau)\equiv d(\Omega(\tau),\Omega(\tau)^{(j)})$. $\Lambda_{\Omega(0)}$ is indeed the network version of a finite, local Lyapunov exponent [44], where $\tau$ is the characteristic time required for the elements inside the $\epsilon$-ball centered at $\Omega(0)$ to diverge to distances of the order of the phase space diameter. For an illustration, in Fig. 2 we depict the distance $d(t)$ between $M=5$ network initializations close to an initial condition $\Omega(0)$ ($\epsilon=10^{-8}$), for a large learning rate $\eta=10$. We observe a clear exponential phase up to $\tau\approx 30$ epochs, the slope denoting the finite local network Lyapunov exponent $\Lambda_{\Omega(0)}$.
Figure 2: Semi-log plot of the evolution (along training) of the network distance $d(t)$ for pairs of network trajectories with closeby initialization $\Omega(0)$, as a function of the number of epochs $t$, for a shallow MLP with $\tanh()$ activation function trained on MNIST with a large learning rate. $d(t)$ displays a stylized exponential expansion followed by saturation. The slope of the exponential phase corresponds to the local network Lyapunov exponent $\Lambda$ and is indicative of chaotic mixing.
Figure 3: Same as Fig 2, but for many different $\epsilon$-balls centered at different initial conditions $\Omega$. Each initial condition leads in principle to a different local network Lyapunov exponent $\Lambda(\Omega)$. In the inset, we display the histogram of local network Lyapunov exponents. The average of this distribution is the estimation of the network MLE $\lambda_{\text{nMLE}}\approx 0.68$.
The network’s Maximum Lyapunov exponent averages the local exponents over different network initializations $\lambda_{\text{nMLE}}=\langle\Lambda_{\Omega(0)}\rangle_{\Omega(0)\in{\cal S}}$. For illustration, this is displayed for $\textsc{Card}({\cal S})=50$ different initial conditions in Fig. 3. The inset of that panel depicts the histogram of local exponents, whose average gives $\lambda_{\text{nMLE}}\approx 0.68$, i.e. the system shows sensitivity to initial conditions. $\lambda_{\text{nMLE}}$ can thus be seen as an order parameter distinguishing two phases: a phase where the search strategy induced by the GD map Eq. 1 is an exploitation one and $\lambda_{\text{nMLE}}\leq 0$, and a phase where the search strategy is of an exploration type with sensitive dependence on initial conditions, and $\lambda_{\text{nMLE}}>0$. The transition between both phases interpolates both search strategies. As a complementary metric that will later assists us in our empirical analysis of the network training dynamics, we also define $\rho$ as the percentage of MLP initializations of the set $\cal S$ for which the $\epsilon$-ball expansion can be well approximated by an exponential function with high statistical significance ($R^{2}>0.9$, $\Lambda>0.05$), i.e this is the percentage of local network Lyapunov exponents which are positive.
Figure 4: Estimation of the network Maximum Lyapunov Exponent $\lambda_{\text{nMLE}}$ for MLP trajectories as a function of the learning rate $\eta$. Error bars denote $\pm$ one standard deviation of the population of finite local network Lyapunov exponents $\{\Lambda(\Omega)\}$. The onset of sensitivity to initial conditions $\lambda_{\text{nMLE}}>0$ marks the change from a purely exploitation-type optimization to an exploration/exploitation type.
Figure 5: Blue diamonds depict $\rho$, the percentage of MLP initializations $\Omega$ leading to training trajectories with positive local Lyapunov exponent $\Lambda(\Omega)>0$ as a function of the learning rate $\eta$. In the same figure, we also plot (red dots) the average training time $\langle\tau\rangle$ (in number of Gradient Descent epochs) needed to reach an accuracy of 90% or larger in the test set. Training is found to be maximally efficient close to the onset of fully-developed sensitivity to initial conditions ($\Lambda(\Omega)>0\ \forall\Omega$).
In Fig. 4 we report $\lambda_{\text{nMLE}}$ as a function of the gradient descent’s learning rate $\eta$. Results indicate that there is a clear transition between exploitation to exploration search, where the exploitation-exploration interpolation hovers in the range $\eta\in[1,10]$. Fig 5 displays $\rho$ as a function of the learning rate. This metric confirms a transition between a phase where no initial conditions display exponential expansion, to a phase where virtually all regions of the phase space display chaotic transients.
Finally, to assess the MLP’s learning and training efficiency in the context of the abovementioned phenomenology, Fig. 5 also depicts the average number of epochs $\langle\tau\rangle$ needed for the MLP to reach an average classification accuracy of at least $0.9$ in the test set (this is averaged over all different MLP initializations in $\cal S$), as a function of the learning rate $\eta$. We find that $\langle\tau\rangle$ displays a non-monotonic shape, and indeed reaches a minimum in the exploitation-exploration interpolation region, at the learning rate $\eta\approx 7.5$, precisely marking the onset of fully-developed sensitivity to initial conditions $\rho\approx 100\%$.
As advanced, this type of optimality in the training dynamics has been recently observed when the largest eigenvalue $\sigma_{\text{max}}$ of the loss function’s Hessian asymptotic converges to $2/\eta$, the so-called edge of stability [41]. In Appendix C Fig. 14, we show how the time series of $\sigma_{\text{max}}(t)$ over training indeed approaches $2/\eta$ when the learning rate hovers around the values for which $\langle\tau\rangle$ is minimized, and suggests that asymptotically self-organizing around the Hessian’s edge-of-stability is precursed by a chaotic transient.
Our results are robust against changes in the classification task (see Appendix B Fig. 8 and 13 for further results on the Iris and CIFAR-10 classification respectively), the type of activation function (see Appendix B Fig. 9 for further results with sigmoid and ReLU functions), the depth of the MLP (see Appendix B Fig. 10 for a comparison of shallow vs deep network), the inclusion of weight regularization (see Appendix B Fig. 11), and also hold for other neural network architectures with inductive biases (see Appendix B Fig. 12 for analysis on a convolutional neural network) and overall highlight the relevance of leveraging transient chaotic mixing in the training of neural networks. In summary, we have found that, as the learning rate increases, the training dynamics transition from a regular, purely exploitation-type dynamics to a chaotic, purely exploration-type dynamics. The transition between both types is rather sharp and occurs in a region that trades-off exploitation and exploration by the onset of a mechanism of chaotic transient –a chaotic fingerprint emerging in the first few dozen epochs of the training dynamics–. This mechanism underpins an efficient search of graph space, eventually leading to faster learning. Evidence suggests that such early-stage, chaotic transient is precursing at a later Hessian trajectories to eventually converge towards their edge of stability.
From a conceptual point of view, our findings suggest a demonstration of Langton’s edge of chaos hypothesis [42], and a confirmation of Verschure’s seminal idea of using chaos as a fast-search mechanism [46]. In hindsight, this is an example where instabilities emerging at the core of numerical schemes –often seen as a nuisance when it comes to e.g. integrating differential equations or root finding– have in the case of searching for neural network representations an unexpected benefit. This benefit occurs because in neural network optimization one uses GD as a search algorithm rather than simply as a local minimizer, and as a consequence, exploitation benefits from exploration when both mechanisms are at play. Aside from this theoretical insight, forcing the training dynamics’ learning rate to be located at the sweet spot could also yield practical benefits. This can be achieved e.g. by using the bisection method to find the sweet spot, by iteratively refining an initial learning rate range $\eta_{\text{min}},\eta_{\text{max}}]$, such that $\rho(\eta_{\text{min}})\approx 0$ and $\rho(\eta_{\text{max}})\approx 100$ as a pre-processing step, before actually training the system.
Finally, further work is needed to clarify the modulating effect of stochastic sources in the network trajectories, such as using stochastic or mini-batch as opposed to full-batch gradient descent [19, 47] or adding dropout, and to address more sophisticated learning rate schedules. Likewise, investigating whether such instability-induced advantage holds in other optimization schemes, or whether other parameters such as the batch size [47] could play the role of a control parameter, and exploring if such phenomenology can be retrieved from low-dimensional projections of network trajectories [48, 49], are all interesting open problems.
Acknowledgments – The authors thank K. Danovski for input in preliminary stages of this project, M. Matías for insightful discussions, A. Fernández Amil for pointing us to [46] after our work was submitted as a preprint and referees for stimulating discussions. PJ ackowledges funding from Maria de Maeztu (MdM) Seal of Excellence (CEX2021-001164-M) via the FPI programme (grant PRE2022-104148), funded by the MICIU/AEI/10.13039/501100011033. LL acknowledges partial support from project CSxAI (PID2024-157526NB-I00) funded by MICIU/AEI/10.13039/501100011033/FEDER, UE. MCS and LL acknowledge partial support from a Maria de Maeztu grant (CEX2021-001164-M) funded by the MICIU/AEI/10.13039/501100011033, and from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL).
Appendix A: Emergence of chaotic dynamics in toy gradient descent maps. Often, gradient descent optimisation of neural networks weights generates monotonically decreasing loss function trajectories towards one of the minima, and common wisdom assumes that such monotonicity is also inherited in weight space. Here we challenge this belief by presenting a toy model for the (gradient descent) optimisation of neural networks which display chaotic trajectories in weight space.
Let us consider a one-dimensional gradient dynamics map
$w_{t+1}=w_{t}-\eta\frac{dV(w)}{dw}\equiv f(w_{t}),$ (A1)
where the learning rate $\eta$ is a map’s parameter and $V(w)$ is a potential function that acts as our simplified loss function. The simplest case of a non-convex loss with multiple minima is perhaps a double well potential
$V(w)=\frac{1}{4}w^{4}-\frac{1}{2}w^{2},$ (A2)
which has two local minima for w=1,1w=-1,1. In this case, $f(w)=w[1-\eta(w^{2}-1)]$ is thus a cubic map. We now show that even in such simple setting, the dynamics can become chaotic for sufficiently large learning rates.
Figure 6: Bifurcation diagram of the one-dimensional gradient map, displaying a period-doubling cascade to chaos.
While for low values of $\eta$ the dynamics of Eq. A1 converge towards one of the two minima, for larger values of this parameter, a bifurcation cascade takes place. In Fig. 6 we plot the bifurcation diagram of the map, obtained numerically by iterating Eq. A1 for 200 steps a random initial conditions $w_{0}\in N(0,0.1)$ (after discarding a transient period) as a function of $\eta$. The figure shows a classical period-doubling bifurcation cascade towards a chaotic state with aperiodic dynamics, intertwined with some self-affine periodic windows. This is the first evidence.
Since this map is one-dimensional, it only has a single Lyapunov exponent $\lambda$, which can be approximated up to arbitrary precision with
$\lambda=\lim_{t\to\infty}\frac{1}{t}\sum_{k=0}^{t-1}|\ln(f^{\prime}(w_{k})|.$ (A3)
In Fig. 7 we show $\lambda$ as a function of the learning rate $\eta$, clearly finding regions for which $\lambda>0$. All in all, this toy model illustrates that gradient descent scheme can yield chaotic dynamics (in the sense of Devaney) for non-convex loss functions (such as the ones involved in neural network optimisation) for sufficiently large learning rates. Since realistic (deep) neural network optimisation just involves substantially more non-convex loss functions of substantially higher-dimensional weight vectors, it is reasonable to expect that moving towards this more complex scenario makes the onset of chaotic behavior for large learning rates a generic phenomenon.
Figure 7: Lyapunov exponent of the gradient map in Eq. A1 as a function of the learning rate $\eta$, showing regions with chaotic behavior.
Appendix B: Onset of sensitivity to initial conditions and training efficiency: Supplementary results. Below we present the results of the key metrics (network Maximum Lyapunov Exponent $\lambda_{\text{nMLE}}$, percentage of positive finite network Lyapunov exponents $\rho$, and average training time $\langle\tau\rangle$) for the additional set of configurations and tasks (Iris vs MNIST vs CIFAR-10, different activation functions, shallow vs deep architecture, effect of regularization and CNNs). These results provide additional evidence for the robustness of the phenomenology described in the main text. Unless otherwise stated, all the networks are initialized with weights drawn from $\mathcal{N}(0,1)$ and biases initialized to zero, and are trained using the same protocol as in the main text. A baseline implementation of the data-generation pipeline used in this work is publicly available at [50].
Figure 8: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$) for each activation function. Results for the Iris task.
We start with the Iris classification task [51] (150 samples, 4 features, 3 classes; 120/30 train/test split). The network architecture is 4–10–3 (83 trainable parameters). Results are shown in Fig. 8. Despite the low dimensionality of the task, the same qualitative transition is observed: the minimum in $\langle\tau\rangle$ coincides with the onset of positive $\lambda_{\mathrm{nMLE}}$.
We next consider the MNIST dataset [43] (60,000/10000 train/test samples; $\times28$ grayscale images) using the 784-64-10 architecture (50,890 parameters). Figure 9 reports results for sigmoid and ReLU activation functions, complementing the tanh case analyzed in the main text. The qualitative structure persists across nonlinearities. For the sigmoid case, the minimum of $\langle\tau\rangle$ extends over a broad range of learning rates, with the subsequent increase appearing only for $\eta\gtrsim 30$.
Figure 9: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$) for the remaining activation functions. Results for the MNIST task.
To assess depth effects, we replace the single hidden layer (64 neurons) by two hidden layers of 32 neurons each (784–32–32–10; 26,506 parameters), keeping total hidden width fixed. Results in Fig. 10 show that the phenomenology persists, although the minimum convergence time increases by approximately a factor of 20 relative to the shallow configuration.
Figure 10: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$). Results for the MNIST task with an extra hidden layer in the architecture and tanh as the activation function.
We further evaluate $L_{2}$-regularized training with loss $\mathcal{L}_{\lambda}=\mathcal{L}+\frac{\lambda}{2}\|\mathscr{W}\|^{2}$ using $\lambda=10^{-3}$ and $10^{-5}$ for the MNIST–tanh configuration. As shown in Fig. 11, the region of optimal training efficiency remains aligned with the onset of positive $\lambda_{\mathrm{nMLE}}$.
Figure 11: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$). Results for the MNIST task with a single hidden layer and tanh activation function, now including $L_{2}$-regularization.
Finally, we consider a convolutional architecture for MNIST consisting of two convolutional layers (8 and 16 kernels, $3\times 3$, padding=1), each follow by $2\times 2$ max pooling, and a fully connected output layer (9,098 trainable parameters). Convolutional layers use ReLU activations and weights are initialized from $\mathcal{N}(0,0.01)$. Training is performed for up to 2,000 epochs with 40 initial conditions and a target accuracy of 80%. As shown in Fig. 12, the characteristic sweet spot in learning rate remains visible.
Figure 12: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$). Results correspond to the CNN applied to the MNIST classification task.
We conclude with the CIFAR-10 dataset [52] (50,000/10,000 train/test samples; $\times32$ RGB images) using the 3072–256–10 MLP architecture (789,258 parameters). Since MLPs typically achieve limited performance on this task, we set the target accuracy to 30%. As shown in Fig. 13, the same qualitative behavior is observed despite the increased task complexity.
Figure 13: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$) for each activation function. Results for the CIFAR-10 task with a target accuracy of 30%.
Appendix C: Evolution of sharpness. Following the approach in [41], we measure the evolution of sharpness during training, defined as the maximum eigenvalue of the Hessian of the training loss with respect to the model parameters. This measure characterises the local curvature of the loss surface and provides insight into the dynamics of the optimisation process. The sharpness is tracked across epochs for the model trained with the MNIST dataset and tanh as activation function.
We use power iteration to approximate the top eigenvalue of the Hessian. This method is computationally efficient and does not require explicit computation and storage of the full Hessian matrix, so it is feasible to apply it during training. We track the evolution of the sharpness over training by measuring the top eigenvalue of the Hessian matrix of the loss function, as shown in Figure 14. For learning rates that minimize $\langle\tau\rangle$, we find that, in line with theoretical considerations [41], sharpness remains bounded once the training has stabilised and closely matches the theoretical limit of $2/\eta$, as indicated by the red dashed line. The results shown correspond to learning rates $\eta=7.5$ and $\eta=10$.
Figure 14: Evolution of sharpness as a function of training epochs. The theoretical value $2/\eta$ is indicated by the dashed red line for reference. Results are shown for learning rates values $\eta=7.5$ and $\eta=10$.